40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1480     Entry time: Tue Apr 14 02:59:02 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Power up until 26 
Yoichi, Peter,

With careful adjustments of the common mode gains, we were able to go up to arm power = 26, sort of robustly (more 50% chance).
At this arm power level, the common mode loop shape still looks good. But the interferometer loses lock easily.
I have to check other DOFs, but the interferometer does not stay locked long enough.
Today, lock losses of the IFO were associated with the lock loss of the PMC whereas the FSS stayed locked.
Probably the AO path got large kicks, which could not be handled by the PMC PZT.

The cause for the IFO lock loss is under investigation.
ELOG V3.1.3-